A block of steel is pulled with constant velocity
Answer #1

I)

pulley 75kg

II)

Applying Newton's second law on both bodies:

block of steel:

N-Wb=0 \Rightarrow N=Wb

T-\mu N=ma=0 since the velocity is constant there is no acceleration

T=\mu W_b (A)

For the weight:

T-W=0\Rightarrow T=W=75\cdot 9.8=735\, N (B)

B in A

\mu =\frac{T}{W_b}=\frac{735}{225\cdot 9.8}=0.333

III)

Now if the rope breaks, T is zero

\mu N=ma\Rightarrow \mu mg=ma\Rightarrow a=\mu g=3.266\, m/s^2

The time i'll take to stop would be:

t=\frac{v_f-v_o}{a}=\frac{0.3}{3.266}=0.091\, s